OpenStudy (anonymous):

Calculate the serie \[\sum_{n=0}^{\infty}(2^{-n}-(\frac{-3}{4})^{n+2})\] Solution \[\sum_{n=0}^{\infty}(2^{-n})-\sum_{n=0}^{\infty}(\frac{-3}{4})^{n+2} = \] \[\sum_{n=0}^{\infty}((\frac{1}{2})^{n})-\sum_{n=0}^{\infty}(\frac{-3}{4})^{n}.(\frac{-3}{4}^{2}) = \] \[\sum_{n=0}^{\infty} \frac{1}{2}^{n} -\sum_{n=0}^{\infty} \frac{-3}{4}^{n}.\sum_{n=0}^{\infty}\frac{9}{16} = \frac{9}{16} \] can you check it out if the solution correct is?

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty}\frac{1}{2^n}=2\] for the first term

OpenStudy (anonymous):

hmm why ? i thought its \[0\]

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty}(\frac{3}{4})^n=\frac{1}{1+\frac{3}{4}}=\frac{7}{4}\] but you have to subtract off the first two terms

OpenStudy (anonymous):

you thought what was zero?

OpenStudy (anonymous):

i thought \[ \sum_{n=0}^{\infty} (2^{-n})\] goes to \[0\]

OpenStudy (anonymous):

ok hold on \[\lim_{n\to \infty}\frac{1}{2^n}=0\] for sure

OpenStudy (anonymous):

otherwise there is no way that the sum would be finite

OpenStudy (anonymous):

but you are adding \[1+\frac{1}{2}+\frac{1}{4}+\frac{1}{8}+\frac{1}{16}+. ..\] and that is clearly not zero, it is 2

OpenStudy (anonymous):

aah ok

OpenStudy (anonymous):

using the formula for the sum of the geometric series \[\sum x^n=\frac{1}{1-x}\] with \(x=\frac[1}{2}\)

OpenStudy (anonymous):

\(x=\frac{1}{2}\)

OpenStudy (anonymous):

that is also how you add up the second series, but you have to be careful, because you are missing the first two terms you are starting at \(n=2\) not \(n=0\) so you sum them using the formula above, subtract off the first two terms, and then was ever you get you subtract from the number \(2\)

OpenStudy (anonymous):

*whatever you get

OpenStudy (anonymous):

sorry something i thought i understand but i didnt why \[1+\frac{1}{2}+\frac{1}{4}+\frac{1}{8}+\frac{1}{16}+. .. = 2 \] ? e.g. why \[\frac{1}{2}+\frac{1}{4}+\frac{1}{8}+\frac{1}{16}+. .. = 1 \] ?

OpenStudy (anonymous):

you want the reason behind the formula, or the formula?

OpenStudy (anonymous):

is there any formula for \[\frac{1}{2}+\frac{1}{4}+\frac{1}{8}+\frac{1}{16}+. .. = 1 \] is ?

OpenStudy (anonymous):

formula for \[a+ar+ar^2+ar^2+...\] is \[\frac{a}{1-r}\]

OpenStudy (anonymous):

lots of ways to see this. for one thing algebra will show you that \[(1-x)(1+x+x^2+x^3+...+x^n)=1-x^{n+1}\] so \[1+x+x^2+x^3+...+x^n=\frac{1-x^{n+1}}{1-x}\] and if \(|x|<1\) we have as \(x\to \infty\) \(x^n\to 0\) leaving \[\sum_{n=0}^{\infty}x^n=\frac{1}{1-x}\]]

OpenStudy (anonymous):

another way is to let \[S=a+ar+ar^2+ar^3+...\] then put \[rS=ar+ar^2+ar^3+...\] so that \[S-rS=a\] and so \[S=\frac{a}{1-r}\]

OpenStudy (anonymous):

hmm ok i try to understand it....

OpenStudy (anonymous):

the problem you posted presupposes you know how to sum a geometric series.

OpenStudy (anonymous):

because to solve the problem you have to sum two of them, then subtract

OpenStudy (anonymous):

hmm ok i look your advices and try to solve it again.. thank you very much

OpenStudy (anonymous):

look over what i wrote because the method is there first sum is 2 second sum i also computed above the only thing left to do is to subtract off the first two terms from second sum, because you are starting at \(n=2\) then whatever you get , subtract that from 2

OpenStudy (anonymous):

that is \[\sum_{n=2}^{\infty}\left(-\frac{3}{4}\right)^n=\sum_{n=0}^{\infty}\left(-\frac{3}{4}\right)^n-(1-\frac{3}{4})\]

OpenStudy (anonymous):

ok i hope i can solve it if not it would be much appreciate if you can solve it for me :)

OpenStudy (anonymous):

i have already more or less \[\sum_{n=2}^{\infty}\left(-\frac{3}{4}\right)^n=\sum_{n=0}^{\infty}\left(-\frac{3}{4}\right)^n-(1-\frac{3}{4})\] \[=\frac{1}{1+\frac{3}{4}}-\frac{1}{4}=\frac{1}{\frac{7}{4}}-\frac{1}{4}=\frac{4}{7}-\frac{1}{4}=\frac{9}{28}\] subtract from 2

OpenStudy (anonymous):

ok is it end solution satellite or i do need subtract it from 2 ?

OpenStudy (anonymous):

aah ok i get it

OpenStudy (anonymous):

yeah you have to subtract it form two

OpenStudy (anonymous):

i need \[2-\frac{9}{28}\]

OpenStudy (anonymous):

thank you very much